OpenStudy (anonymous):

Given f(x) = x^2/2-x, find f(1-i√2) Please Help

OpenStudy (anonymous):

I tried so many ways but I just can't get it Please help

OpenStudy (anonymous):

There are 4 choices It could be A. -3-2i√2 B. (-5-i√2)/3 C. (-7-5i√2)/3 D. (-7+5i√)/3

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

f(x) = x^2/2-x, find f(1-i√2) \[f(x)= \frac{ x ^{2} }{ 2-x }\] \[i.e. f(1-i \sqrt{2})= \frac{ (1-i \sqrt{2}) ^{2} }{ 2-(1-i \sqrt{2}) }\] \[=\frac{ 1+2i ^{2} -2i \sqrt{2}}{ 2-1+i \sqrt{2} }= \frac{1-2 -2i \sqrt{2}}{ 1+i \sqrt{2} }\] \[= \frac{-1 -2i \sqrt{2}}{ 1+i \sqrt{2} }= -\frac{(1+ 2i \sqrt{2})}{ (1+i \sqrt{2}) }\] (Now we will rationalize the denominator) \[=-\frac{(1+ 2i \sqrt{2})}{ (1+i \sqrt{2}) }\times \frac{ (1-i \sqrt{2}) }{ (1-i \sqrt{2}) }\] \[=-\frac{1+ 2i \sqrt{2}-i \sqrt{2}-4i ^{2}}{ (1-2i ^{2}) }\] =\[=-\frac{1+ i \sqrt{2}+4}{ 1+2 } =-\frac{5+ i \sqrt{2}}{3}\] and this is the solution to the given problem.

OpenStudy (anonymous):

Thus the option B is correct.

OpenStudy (anonymous):

I am really sorry but I don't understand this at all. Can you put it in simpler terms

OpenStudy (anonymous):

There is no simpler way to understand this than this method. first of all u need to be aware with concept of complex numbers i.e. i^2 = -1. Except this general calculations has been used to solve it

OpenStudy (anonymous):

In the given problem i have substituted the given value of x.

OpenStudy (anonymous):

I tried that but I got a different answer then the choices above.

OpenStudy (anonymous):

Follow the method i have posted here, you will definitely reach to your result.

OpenStudy (anonymous):

Thank you so much